# Time-Varying Gaussian Process Bandits with Unknown Prior

### v3

@@ 5 PROOF OF THE REGRET BOUND
where $\xi_{t}=2R^{2}\log\frac{|\mathcal{U}|\pi^{2}t^{2}}{6\delta_{B}}$.
@@
For $f\sim p(f)$ with some mean and kernel functions $\mu^{p}(\cdot)$ and $k^{p}(\cdot,\cdot)$ with properties $|\mu^{p}(\cdot)|<\infty$ and $k^{p}(\cdot,\cdot)\leq 1$, there exists a constant $B>0$ s.t.:
$\max_{t\in[T]}\sup_{\bm{x}\in\mathcal{X}}\left|f(\bm{x},t)\right|\leq B$
as long as $|\mathcal{X}|<\infty$ and the event of Theorem 3.1 holds or $|\mathcal{X}|=\infty$ and event of Theorem 3.3 holds.
Observe that by union bound Theorem 3.1 and Lemma 5.1 (and thus also Lemma 5.2) hold together with probability at least $1-\delta_{A}-\delta_{B}$. We choose $\delta_{A}=\delta_{B}=\frac{\delta}{2}$. We prove the regret bound, assuming the probabilistic statements in Theorem 3.1 and Lemma 5.1 hold, and as such the resulting bound holds with probability at least $1-\delta$.
@@
$=\left|\sum_{i\in S^{p^{*}}_{t}}\left(y_{i}-f(\bm{x}_{i},i)+f(\bm{x}_{i},i)-\mu^{p^{*}}_{i-1}(\bm{x}_{i},i)\right)\right|$
$\leq\left|\sum_{i\in S^{p^{*}}_{t}}\epsilon_{i}\right|+\sum_{i\in S^{p^{*}}_{t}}\left|f(\bm{x}_{i},i)-\mu^{p^{*}}_{i-1}(\bm{x}_{i},i)\right|$
$\leq\sqrt{\xi_{t}|S^{p^{*}}_{t}|}+\sum_{i\in S^{p^{*}}_{t}}\beta_{i}^{p^{*}}\sigma_{i}^{p^{*}}(\bm{x}_{i},i),$
@@
$f(\bm{x}^{*}_{t},t)$ $\leq\mu_{t-1}^{p^{*}}(\bm{x}^{*}_{t},t)+\beta_{t}^{p^{*}}\sigma^{p^{*}}_{t-1}(\bm{x}^{*}_{t},t)$ (1)
$\leq\max_{\bm{x}_{t}\in\mathcal{X}_{t}}\max_{p\in\mathcal{U}_{t}}\mu_{t-1}^{p}(\bm{x}_{t},t)+\beta_{t}^{p}\sigma_{t-1}^{p}(\bm{x}_{t},t),$ (2)
where first inequality is due to Theorem 3.1. For simplicity we will write $\mu_{t-1}^{+}=\mu_{t-1}^{p_{t}}$, $\sigma_{t-1}^{+}=\sigma_{t-1}^{p_{t}}$ and $\beta^{+}_{t}=\beta_{t}^{p_{t}}$. Having established an upper bound on $f(\bm{x}^{*}_{t},t)$, it remains to find a lower bound on $f(\bm{x}_{t},t)$. We show the following:
$f(\bm{x}_{t},t)\geq\mu^{+}_{t-1}(\bm{x}_{t},t)-\eta_{t}+\epsilon_{t}.$ (3)
The observation above is a key component of our proof, as it relates the lower bound on the function to the prediction error of $\eta_{t}$ of the model defined by prior $p_{t}$. Combining Inequalities (2)-(3) yields the following bound on instantaneous regret:
$r_{t}\leq\beta_{t}^{+}\sigma^{+}_{t-1}(\bm{x}_{t},t)-\eta_{t}+\epsilon_{t}.$ (4)
@@
$\mathcal{C}=\left\{t\in[T]:\left|\sum_{i\in S^{p_{t}}}\eta_{i}\right|>\sqrt{\xi_{t}|S^{p_{t}}|}+\sum_{i\in S^{p_{i}}}\beta^{p_{i}}_{i}\sigma_{i}^{p_{t}}(\bm{x}_{i},i)\right\}$
Notice that for each $t\in\mathcal{C}$, we discard one possible prior and as such $|\mathcal{C}|\leq|\mathcal{U}|$. On those iterations, we may suffer an arbitrarily high regret. Due to Lemma 5.2 we have $r_{t}\leq 2B$ for all $t\in\mathcal{C}$. Using Equation 4, we thus see that the cumulative regret is bounded as:
@@
$\leq 2|\mathcal{U}|B+\sum_{t\notin\mathcal{C}}\beta^{+}_{t}\sigma^{+}_{t-1}(\bm{x}_{t},t)+\sum_{p\in\mathcal{U}}\sum_{t\in S^{u}_{T}\setminus\mathcal{C}}\left(\epsilon_{t}-\eta_{t}\right).$
@@
$\sum_{p\in\mathcal{U}}\sum_{t\in S^{u}_{T}\setminus\mathcal{C}}-\eta_{t}\leq\sum_{p\in\mathcal{U}}\sqrt{\xi_{T}|S^{p}_{T}|}+\sum_{p\in\mathcal{U}}\sum_{t\in S^{u}_{T}\setminus\mathcal{C}}\beta^{p}_{t}\sigma_{t-1}^{p}(\bm{x}_{t},t).$
@@
$R_{T}$ $\leq 2|\mathcal{U}|B+2\sum_{p\in\mathcal{U}}\sqrt{\xi_{T}|S^{p}_{T}|}+2\sum_{t\notin\mathcal{C}}\beta_{t}^{+}\sigma^{+}_{t-1}(\bm{x}_{t},t)$
$\leq 2|\mathcal{U}|B+2\sqrt{\xi_{T}|\mathcal{U}|T}+2\sum_{t\notin\mathcal{C}}\beta_{t}^{+}\sigma^{+}_{t-1}(\bm{x}_{t},t),$
@@
$\sum_{t\notin\mathcal{C}}\beta^{+}_{t}\sigma^{+}_{t-1}(\bm{x}_{t},t)\leq\beta_{T}\sqrt{CT|\mathcal{U}|\gamma_{T}},$
@@
$R_{T}\leq 2|\mathcal{U}|B+2\sqrt{|\mathcal{U}|\xi_{T}T}+2\beta_{T}\sqrt{CT|\mathcal{U}|\gamma_{T}},$
@@ 6 DISCUSSION
The first term in our bound $|\mathcal{U}|B$ does not depend on time and as such its impact on the bound is negligible for sufficiently large $T$. As for most typically used kernel (RBF, Matern), the MIG $\gamma_{T}$ is at least logarithmic in $T$ (Srinivas et al., 2010), the third term dominates the second term. In case where prior is known, GP-UCB (and its time-varying variant) admits a regret bound of $R_{T}\leq\mathcal{O}(\beta_{T}\sqrt{T\gamma_{T}^{p^{\star}}})$ (Srinivas et al., 2010; Bogunovic et al., 2016). As such, our bound is $\sqrt{|U|}$ factor away and scales with $\gamma_{T}=\max_{p\in\mathcal{U}}\gamma_{T}^{p}$ instead of $\gamma^{p^{\star}}_{T}$. However, in many applications, $\gamma^{p}_{T}$ will be the same for all priors $p\in\mathcal{U}$, for example, in the case where priors differ with respect to mean only, and then $\gamma_{T}=\gamma_{T}^{p^{\star}}$. Even if this is not the case, suffering regret scaling with $\gamma_{T}$ is still much better than the arbitrarily high regret we may suffer if we rely on the wrong prior, which might happen if the prior is chosen by an alternative method without a correctness guarantee.
@@ 7 EXPERIMENTS
We now empirically evaluate our proposed algorithm. We conduct the evaluation on both toy and real-world problems. The first baseline we compare against is the MLE, which simply selects the prior that produces a model with the highest marginal likelihood, i.e. $p_{t}=\arg\max_{p\in\mathcal{U}}P(\mathcal{D}|p)$ and then optimises the UCB function $\textrm{UCB}_{t}^{p_{t}}(\bm{x},t)$ produced by that model. Another baseline is Fully Bayesian treatment of the prior, which maximises the weighted average of UCB functions obtained under different priors with the weights being the marginal likelihood of the model with given prior, i.e. $\sum_{p\in\mathcal{U}}\textrm{UCB}_{t}^{p}(\bm{x},t)P(p|\mathcal{D}_{t-1})$, where $P(p|\mathcal{D}_{t-1})\propto P(\mathcal{D}_{t-1}|p)P(p)$ and $P(p)$ is the hyperprior (that is prior over priors). In our experiments we simply take $P(p)=\frac{1}{|\mathcal{U}|}$. The final baseline we compare against is Regret Balancing, which uses the observed function values in selecting the priors the use at future iterations. For completeness, we provide a pseudo-code in Supplementary Material Section D of the exact version of Regret Balancing that we used. Such an algorithm admits a regret bound, but only in the case, where the problem is stationary. As such, comparing the performance of Regret Balancing to PE-GP-UCB will help gauge the improvement coming from not relying on the stationarity of the objective. We detail each experiment below. To run all experiments we use one machine with AMD Ryzen 7 5800X 8-Core processor, no GPU was required.
@@
Toy Problem As the first problem, we consider a toy set, where the ground truth function $f$ is a sample from true GP. The function $f$ is stationary, i.e. $\forall_{t,t^{\prime}\in[T]}f(\cdot,t)=f(\cdot,t^{\prime})$ for simplicity. The true GP employs an RBF kernel with a specially designed mean function that features ten distinct ‘hills’—regions where the mean function is higher. We use prior 2 from Figure 3 as the true mean function, which has one prominent hill around 0.1 and nine smaller hills. For the estimation setup, we prepare eleven GP prior candidates, each using the same RBF kernel but with different mean functions, as shown in Figure 3. Each prior features a tall hill at varying locations, along with nine smaller hills. Prior 0, however, only has ten short hills. An algorithm that can correctly identify prior 2 should achieve faster regret convergence. Furthermore, we introduce a time-varying search space $\mathcal{X}_{t}$, which alternates between the entire domain $\mathcal{X}$ and a restricted search space that excludes the true tall hill, creating an additional challenge for estimation (see our code for more details).
@@
We plot the cumulative regret of each algorithm in Figure 2. We can see that MLE, Fully Bayesian approach, and Regret Balancing all perform similarly and outperform Random Search, however, still underperform compared to our algorithm PE-GP-UCB. The histogram in Figure 2 sheds some light as to why this is happening. MLE keeps selecting the same (wrong prior) and gets stuck at suboptimal values. Regret Balancing is trying out all of the priors in a roughly uniform manner, however, since the true prior is selected only a fraction of the time and half of the time the optimal region is excluded from the search space, it is likely that the algorithm will not observe better function values while using the true prior. As a result, the Regret Balancing scheme will not commit to the true prior and keep exploring indefinitely. On the other hand, our algorithm PE-GP-UCB, explores all the priors and swiftly rejects incorrect ones, committing to the true prior, which results in the lowest regret.
@@
Intel Temperature Dataset As the second experiment, we consider a real-world dynamic optimisation on the famous Intel Research Dataset [^1] https://db.csail.mit.edu/labdata/labdata.html. The dataset contains temperature measurements gathered from over 50 sensors located in the Intel Research office in Berkeley. The goal is to select the sensor with the highest temperature at each timestep and the regret is the difference between the highest and selected temperature. For the sake of optimisation, we use the data coming from 14th March 2004 and we define each timestep to be a ten-minute interval. To obtain the prior of the function, we use the previous days in March, obtaining a total of thirteen priors (one from each day). Following previous work Bogunovic et al. (2016), for each prior we use a zero mean function and a kernel of form $k((i,t),(i^{\prime},t^{\prime}))=\textrm{Cov}(i,i^{\prime})(1-\epsilon)^{|t-t^{\prime}|/2}$, where $\textrm{Cov}(i,i^{\prime})$ is the empirical covariance between sensors $i$ and $i^{\prime}$ (as estimated on a given day) and $\epsilon$ is selected to produce the best fit on that day. Since we do not know which day would be most similar to the 14th of March, we do not know which prior we should rely on. We run our method and the baseline and show the plot of average regret against time step in Figure 4.

### v4

@@ 5 PROOF OF THE REGRET BOUND
where $\xi_{t}=2R^{2}\log\frac{|\mathcal{U}|\pi^{2}t^{2}}{3\delta_{B}}$.
@@
For $f\sim p(f)$ with some mean and kernel functions $\mu^{p}(\cdot)$ and $k^{p}(\cdot,\cdot)$ with properties $|\mu^{p}(\cdot)|<\infty$ and $k^{p}(\cdot,\cdot)\leq 1$, for any $t\leq T$ with probability at least $1-\delta_{C}$ we have:
$\max_{t^{\prime}\in[t]}\sup_{\mathbf{x}\in\mathcal{X}}\left|f(\mathbf{x},t^{\prime})\right|\leq B_{t},$
where $B_{t}=\beta_{t}^{p^{\star}}+\max_{t^{\prime}\in[t]}\sup_{\mathbf{x}\in\mathcal{X}}|\mu^{p^{\star}}(\mathbf{x},t^{\prime})|+1$ and $\beta_{T}^{p^{\star}}$ is to be set as in Theorem 3.1 for discrete case and as in Theorem 3.3 for continuous case with $\delta_{A}$ replaced with $\delta_{C}$.
Observe that by union bound Theorem 3.1 and Lemma 5.1 and Lemma 5.2) hold together with probability at least $1-\delta_{A}-\delta_{B}-\delta_{C}$. We choose $\delta_{A}=\delta_{B}=\delta_{C}=\frac{\delta}{3}$. We prove the regret bound, assuming the probabilistic statements in Theorem 3.1 and Lemma 5.1 hold, and as such the resulting bound holds with probability at least $1-\delta$.
@@
$=\left|\sum_{i\in S^{p^{*}}_{t}}\left(y_{i}-f(\mathbf{x}_{i},i)+f(\mathbf{x}_{i},i)-\mu^{p^{*}}_{i-1}(\mathbf{x}_{i},i)\right)\right|$
$\leq\left|\sum_{i\in S^{p^{*}}_{t}}\epsilon_{i}\right|+\sum_{i\in S^{p^{*}}_{t}}\left|f(\mathbf{x}_{i},i)-\mu^{p^{*}}_{i-1}(\mathbf{x}_{i},i)\right|$
$\leq\sqrt{\xi_{t}|S^{p^{*}}_{t}|}+\sum_{i\in S^{p^{*}}_{t}}\beta_{i}^{p^{*}}\sigma_{i}^{p^{*}}(\mathbf{x}_{i},i),$
@@
$f(\mathbf{x}^{*}_{t},t)$ $\leq\mu_{t-1}^{p^{*}}(\mathbf{x}^{*}_{t},t)+\beta_{t}^{p^{*}}\sigma^{p^{*}}_{t-1}(\mathbf{x}^{*}_{t},t)$ (1)
$\leq\max_{\mathbf{x}_{t}\in\mathcal{X}_{t}}\max_{p\in\mathcal{U}_{t}}\mu_{t-1}^{p}(\mathbf{x}_{t},t)+\beta_{t}^{p}\sigma_{t-1}^{p}(\mathbf{x}_{t},t),$ (2)
where first inequality is due to Theorem 3.1. For simplicity we will write $\mu_{t-1}^{+}=\mu_{t-1}^{p_{t}}$, $\sigma_{t-1}^{+}=\sigma_{t-1}^{p_{t}}$ and $\beta^{+}_{t}=\beta_{t}^{p_{t}}$. Having established an upper bound on $f(\mathbf{x}^{*}_{t},t)$, it remains to find a lower bound on $f(\mathbf{x}_{t},t)$. We have also that:
$f(\mathbf{x}_{t},t)=y_{t}-\epsilon_{t}=\mu^{+}_{t-1}(\mathbf{x}_{t},t)+\eta_{t}-\epsilon_{t}.$ (3)
The observation above is a key component of our proof, as it relates the function value to the prediction error of $\eta_{t}$ of the model defined by prior $p_{t}$. Combining (2)-(3) yields the following bound on instantaneous regret:
$r_{t}=f(\mathbf{x}^{*}_{t},t)-f(\mathbf{x}_{t},t)\leq\beta_{t}^{+}\sigma^{+}_{t-1}(\mathbf{x}_{t},t)-\eta_{t}+\epsilon_{t}.$ (4)
@@
$\mathcal{C}=\left\{t\in[T]:\left|\sum_{i\in S^{p_{t}}}\eta_{i}\right|>\sqrt{\xi_{t}|S^{p_{t}}|}+\sum_{i\in S^{p_{i}}}\beta^{p_{i}}_{i}\sigma_{i}^{p_{t}}(\mathbf{x}_{i},i)\right\}$
Notice that for each $t\in\mathcal{C}$, we discard one possible prior and as such $|\mathcal{C}|\leq|\mathcal{U}|$. On those iterations, we may suffer an arbitrarily high regret. Due to Lemma 5.2 we have $r_{t}\leq 2B_{t}$ for all $t\in\mathcal{C}$. At each timestep, we can reject a prior and expand $\mathcal{C}$ or not reject any of the priors and then $\mathcal{C}$ does not grow. As there is finite number of priors we must have that $|\mathcal{C}|<|\mathcal{U}|$ and there must be some final iteration $t^{\star}$ after which $\mathcal{C}$ was never expanded anymore. As $B_{t}$ can only grow with $t$, we have that $\forall_{t\in\mathcal{C}}B_{t}\leq B_{t^{\star}}$ and thus $\forall_{t\in\mathcal{C}}r_{t}\leq 2B_{t^{\star}}$. Using Equation 4, we thus see that the cumulative regret is bounded as:
@@
$\leq 2|\mathcal{U}|B_{t^{\star}}+\sum_{t\notin\mathcal{C}}\beta^{+}_{t}\sigma^{+}_{t-1}(\mathbf{x}_{t},t)+\sum_{p\in\mathcal{U}}\sum_{t\in S^{u}_{T}\setminus\mathcal{C}}\left(\epsilon_{t}-\eta_{t}\right).$
@@
$\sum_{p\in\mathcal{U}}\sum_{t\in S^{u}_{T}\setminus\mathcal{C}}-\eta_{t}\leq\sum_{p\in\mathcal{U}}\sqrt{\xi_{T}|S^{p}_{T}|}+\sum_{p\in\mathcal{U}}\sum_{t\in S^{u}_{T}\setminus\mathcal{C}}\beta^{p}_{t}\sigma_{t-1}^{p}(\mathbf{x}_{t},t).$
@@
$R_{T}$ $\leq 2|\mathcal{U}|B_{t^{\star}}+2\sum_{p\in\mathcal{U}}\sqrt{\xi_{T}|S^{p}_{T}|}+2\sum_{t\notin\mathcal{C}}\beta_{t}^{+}\sigma^{+}_{t-1}(\mathbf{x}_{t},t)$
$\leq 2|\mathcal{U}|B_{t^{\star}}+2\sqrt{\xi_{T}|\mathcal{U}|T}+2\sum_{t\notin\mathcal{C}}\beta_{t}^{+}\sigma^{+}_{t-1}(\mathbf{x}_{t},t),$
@@
$\sum_{t\notin\mathcal{C}}\beta^{+}_{t}\sigma^{+}_{t-1}(\mathbf{x}_{t},t)\leq\beta_{T}\sqrt{CT|\mathcal{U}|\gamma_{T}},$
@@
$R_{T}\leq 2|\mathcal{U}|B_{t^{\star}}+2\sqrt{\xi_{T}|\mathcal{U}|T}+2\beta_{T}\sqrt{CT|\mathcal{U}|\gamma_{T}},$
@@ 6 DISCUSSION
The first term in our bound $|\mathcal{U}|B_{t^{\star}}$ depends on $B_{t^{\star}}=\beta_{1}+\max_{t^{\prime}\in[t^{\star}]}\sup_{\mathbf{x}\in\mathcal{X}}|\mu^{p^{\star}}(\mathbf{x},t^{\prime})|+1$. Note that as there is a finite number of priors, we cannot reject priors infinitely, as stated by next Lemma.
Lemma 6.1.
We have that $P(\lim_{T\to\infty}\frac{t^{\star}}{T}\to 0)=1$ and $P(\lim_{T\to\infty}\frac{B_{t^{\star}}}{T}\to 0)=1$.
As such for sufficiently large $T$, $t^{\star}$ will not depend on $T$ and as such $B_{t^{\star}}$ will also not depend on $T$ and the first term will get dominated by the remaining terms. As for most commonly used kernels, the MIG $\gamma_{T}$ is at least logarithmic in $T$ (Srinivas et al., 2010), the third term dominates the second term. In case where prior is known, GP-UCB (and its time-varying variant) admits a regret bound of $R_{T}\leq\mathcal{O}(\beta_{T}\sqrt{T\gamma_{T}^{p^{\star}}})$ (Srinivas et al., 2010; Bogunovic et al., 2016). As such, our bound is $\sqrt{|U|}$ factor away and scales with $\gamma_{T}=\max_{p\in\mathcal{U}}\gamma_{T}^{p}$ instead of $\gamma^{p^{\star}}_{T}$. This is an artifact of the fact that PE-GP-UCB is doubly-optimistic and if one prior produces upper confidence bound that is always the highest among other priors, it will always be selected as $p_{t}$. As such, PE-GP-UCB cannot be used when priors differ only with respect to outputscale. However, in many applications, $\gamma^{p}_{T}$ will be the same for all priors $p\in\mathcal{U}$, for example, in the case where priors differ with respect to mean only, and then $\gamma_{T}=\gamma_{T}^{p^{\star}}$. Even if this is not the case, suffering regret scaling with $\gamma_{T}$ is still much better than the arbitrarily high regret we may suffer if we rely on the wrong prior, which might happen if the prior is chosen by an alternative method without a correctness guarantee.
The scaling with the square root of the number of possible priors $\sqrt{|U|}$ can be problematic when $|U|$ is large, but this rate is actually better than for Regret Balancing (Pacchiano et al., 2020; Abbasi-Yadkori et al., 2020), where the time-dependent part of bound grows linearly with number of possible models. In general, linear dependence on the number of base algorithms is common among Master Algorithms literature, where, for example, the celebrated CORRAL algorithm (Agarwal et al., 2017) also achieves at least linear scaling. As such, we believe the square root scaling $\sqrt{|U|}$ is relatively favourable, when compared to algorithms solving similar problems.
Our bound can be applied to different types of kernels, given that their MIG is known. For concreteness, we now plug the expression from (Bogunovic et al., 2016) for MIG of two commonly used time-varying kernels.
Corollary 6.2.
Under the same conditions as Theorem 4.1, the following statements hold:
  • Let all candidate priors utilise the squared exponential kernel, we then have that regret of PE-GP-UCB: $R_{T}=\tilde{\mathcal{O}}(|\mathcal{U}|B_{t^{\star}}+\sqrt{|\mathcal{U}|}\max\{T\varepsilon^{1/6},\sqrt{T}\})$
  • Let all candidate priors utilise the $\nu$-Matern kernel, we then have that regret of PE-GP-UCB: $R_{T}=\tilde{\mathcal{O}}(|\mathcal{U}|B_{t^{\star}}+\sqrt{|\mathcal{U}|}\max\{T\varepsilon^{\frac{1}{2}\frac{1-c}{3-c}},\sqrt{T^{1+c}}\})$
@@
On high level the prior rejection mechanism might resemble the soft revision mechanism (Cattaneo, 2014; Augustin & Schollmeyer, 2021). However, there is a number of important differences. Soft-revision operates on the likelihood, whereas we operate on the predictive error of the models. While linking the likelihood-based approach to guarantees on cumulative regret bounds may not be straightforward, in our proof we are able to express the regret directly in terms of the predictive error, on which our algorithm operates, allowing us to link these two notions. Soft revisions also rely on finding the prior producing the biggest marginal to decide which priors to reject, whereas in our approach the rejection is done regardless of how well other priors perform.
@@ 7 EXPERIMENTS
We now empirically evaluate our proposed algorithm. We conduct the evaluation on both toy and real-world problems. The first baseline we compare against is the MLE, which simply selects the prior that produces a model with the highest marginal likelihood, i.e. $p_{t}=\arg\max_{p\in\mathcal{U}}P(\mathcal{D}|p)$ and then optimises the UCB function $\textrm{UCB}_{t}^{p_{t}}(\mathbf{x},t)$ produced by that model. Another baseline is Fully Bayesian treatment of the prior, which maximises the weighted average of UCB functions obtained under different priors with the weights being the marginal likelihood of the model with given prior, i.e. $\sum_{p\in\mathcal{U}}\textrm{UCB}_{t}^{p}(\mathbf{x},t)P(p|\mathcal{D}_{t-1})$, where $P(p|\mathcal{D}_{t-1})\propto P(\mathcal{D}_{t-1}|p)P(p)$ and $P(p)$ is the hyperprior (that is prior over priors). In our experiments we simply take $P(p)=\frac{1}{|\mathcal{U}|}$. The final baseline we compare against is Regret Balancing, which uses the observed function values in selecting the priors the use at future iterations. For completeness, we provide a pseudo-code in Supplementary Material Section D of the exact version of Regret Balancing that we used. Such an algorithm admits a regret bound, but only in the case, where the problem is stationary. As such, comparing the performance of Regret Balancing to PE-GP-UCB will help gauge the improvement coming from not relying on the stationarity of the objective. We detail each experiment below. Our code is available on github [^1] https://github.com/JuliuszZiomek/PE-GP-UCB. To run all experiments we use one machine with AMD Ryzen 7 5800X 8-Core processor, no GPU was required.
@@
Toy Problem As the first problem, we consider a toy setup, where the ground truth function $f$ is a sample from a GP. The function $f$ is itself stationary, i.e. $\forall_{t,t^{\prime}\in[T]}f(\cdot,t)=f(\cdot,t^{\prime})$, but the space of available points to query changes between iterations, making the problem non-stationary. The true GP employs an RBF kernel with a specially designed mean function that features ten distinct ‘hills’—regions where the mean function is higher. We use prior 2 from Figure 3 as the true mean function, which has one prominent hill around 0.1 and nine smaller hills. For the estimation setup, we prepare eleven GP prior candidates, each using the same RBF kernel but with different mean functions, as shown in Figure 3. Each prior features a tall hill at varying locations, along with nine smaller hills. Prior 0, however, only has ten short hills. An algorithm that can correctly identify prior 2 should achieve faster regret convergence. Furthermore, we introduce a time-varying search space $\mathcal{X}_{t}$, which alternates between the entire domain $\mathcal{X}$ and a restricted search space that excludes the true tall hill, creating an additional challenge for estimation. As such, the optimal region is excluded from the available query points at every even iteration.
@@
We plot the cumulative regret of each algorithm in Figure 2. We can see that MLE, Fully Bayesian approach, and Regret Balancing all perform similarly and outperform Random Search, however, still underperform compared to our algorithm PE-GP-UCB. The histogram in Figure 2 sheds some light as to why this is happening. MLE keeps selecting the same (wrong prior) and gets stuck at suboptimal values. Regret Balancing is trying out all of the priors in a roughly uniform manner, however, since the true prior is selected only a fraction of the time and half of the time the optimal region is excluded from the search space, it is likely that the algorithm will not observe better function values while using the true prior. As a result, the Regret Balancing scheme will not commit to the true prior and keep exploring indefinitely. On the other hand, our algorithm PE-GP-UCB, explores all the priors and swiftly rejects incorrect ones, committing to the true prior, resulting in the lowest regret.
@@
Intel Temperature Dataset As the second experiment, we consider a real-world dynamic optimisation on the famous Intel Research Dataset [^2] https://db.csail.mit.edu/labdata/labdata.html. The dataset contains temperature measurements gathered from over 50 sensors located in the Intel Research office in Berkeley. The goal is to select the sensor with the highest temperature at each timestep and the regret is the difference between the highest and selected temperature. For the sake of optimisation, we use the data coming from 14th March 2004 and we define each timestep to be a ten-minute interval. To obtain the prior of the function, we use the previous days in March, obtaining a total of thirteen priors (one from each day). Following previous work Bogunovic et al. (2016), for each prior we use a zero mean function and a kernel of form $k((i,t),(i^{\prime},t^{\prime}))=\textrm{Cov}(i,i^{\prime})(1-\varepsilon)^{|t-t^{\prime}|/2}$, where $\textrm{Cov}(i,i^{\prime})$ is the empirical covariance between sensors $i$ and $i^{\prime}$ (as estimated on a given day) and $\varepsilon$ is selected to produce the best fit on that day. Since we do not know which day would be most similar to the 14th of March, we do not know which prior we should rely on. We run our method and the baseline and show the plot of average regret against time step in Figure 4.